OpenStudy (anonymous):

At noon, ship A is 50km west of ship B. Ship A is sailing south at 30km/h and ship B is sailing north at 20km/h. How fast is the distance between the ships changing at 4:00pm.

OpenStudy (anonymous):

this is what i did but the answer is wrong

OpenStudy (anonymous):



OpenStudy (anonymous):

yes i think i got the same diagram

OpenStudy (anonymous):

yes, i agree

OpenStudy (anonymous):

did i do it a different way in the picture above?

OpenStudy (anonymous):

so the change in d is (1/2)(50^2t)/(sqrt((50t)^2 +50^2) )

OpenStudy (anonymous):

left out a square on the 50... fixed it..

OpenStudy (anonymous):

so whats t?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

let me try it 1sec

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } \frac{ 50^2t }{ \sqrt{(50t)^2 +50^2} }\]

OpenStudy (anonymous):

~24.25

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

i tried again and got 24.25 and its wrong

OpenStudy (anonymous):

dang... yeah, I just caught the error

OpenStudy (anonymous):

there's no 1/2 out front because t^2 ' =2t

OpenStudy (anonymous):

\[\frac{ 50^2*t }{\sqrt{(50t)^2 +50^2 } }\]

OpenStudy (anonymous):

so 48.5

OpenStudy (anonymous):

got it. perfect, thanks

OpenStudy (anonymous):

sure:)